# A Comprehensive Survey of Dataset Distillation

### v2

@@ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$
FTD [29] - - - - - - $\mathbf{73.8}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ $\mathbf{65.7}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ -
TESLA [25] - - - $\mathbf{66.4}$ $\scriptscriptstyle\mathbf{\pm 0.8}$ $\mathbf{48.9}$ $\scriptscriptstyle\mathbf{\pm 2.2}$ - - - -
@@
CAFE [113] - - - - - -
$55.5$
$\scriptscriptstyle\pm 0.4$
$25.3$
$\scriptscriptstyle\pm 0.9$
KIP [87] $\mathbf{49.9}$ $\scriptscriptstyle\mathbf{\pm 0.2}$
@@
$\mathbf{16.6}$ $\scriptscriptstyle\mathbf{\pm 1.4}$
@@
Table 4: Cross-architecture performance w.r.t. ResNet-10 (RN-10) and DenseNet-121 (DN-121) on CIFAR-10. The results are derived from Lee et al. [62]
Methods IPC=11 IPC=$10$ IPC=$50$
ConvNet RN-10 DN-121 ConvNet RN-10 RN-121 ConvNet RN-10 RN-121
DSA [126]
$28.8$
$\scriptscriptstyle\pm 0.7$
$25.1$
$\scriptscriptstyle\pm 0.8$
$25.9$
$\scriptscriptstyle\pm 1.8$
$52.1$
$\scriptscriptstyle\pm 0.5$
$31.4$
$\scriptscriptstyle\pm 0.9$
$32.9$
$\scriptscriptstyle\pm 1.0$
$60.6$
$\scriptscriptstyle\pm 0.5$
$49.0$
$\scriptscriptstyle\pm 0.7$
$53.4$
$\scriptscriptstyle\pm 0.8$
DM [128]
$26.0$
$\scriptscriptstyle\pm 0.8$
$13.7$
$\scriptscriptstyle\pm 1.6$
$12.9$
$\scriptscriptstyle\pm 1.8$
$48.9$
$\scriptscriptstyle\pm 0.6$
$31.7$
$\scriptscriptstyle\pm 1.1$
$32.2$
$\scriptscriptstyle\pm 0.8$
$63.0$
$\scriptscriptstyle\pm 0.4$
$49.1$
$\scriptscriptstyle\pm 0.7$
$53.7$
$\scriptscriptstyle\pm 0.7$
@@
KFS [62] $\mathbf{59.8}$ $\scriptscriptstyle\mathbf{\pm 0.5}$ $\mathbf{47.0}$ $\scriptscriptstyle\mathbf{\pm 0.8}$ $\mathbf{49.5}$ $\scriptscriptstyle\mathbf{\pm 1.3}$ $\mathbf{72.0}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{70.3}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{71.4}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ $\mathbf{75.0}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ $\mathbf{75.1}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{76.3}$ $\scriptscriptstyle\mathbf{\pm 0.4}$
@@ 6.1 Standard Benchmark
For a comprehensive comparison, we also present the test accuracy w.r.t. a random select search, coreset approach, and the original target dataset. For the coreset approach, the Herding algorithm [94] is employed in this survey due to its superior performance. Notably, for most of parametric DD methods, ConvNet [33] is the default architecture to distill the synthetic data; and all the test accuracy of distilled data are obtained by training the same architecture of ConvNet for fair comparison.
The empirical comparison results associated with MNIST, FashionMNIST, SVHN, CIFAR-10/100, and Tiny ImageNet are presented in Table 1. In spite of severe scale issues and few implementations on ImageNet-1K, we present the DD results of ImageNet-1K in Table 2 for completion. As shown in these tables, many methods are not tested on some datasets due to the lack of meaning for easy datasets or scalability problems for large datasets. We preserve these blanks in tables for a more fair and comprehensive comparison. To make the comparison clear, we use the orange number for the best of two nonfactorized DD methods and blue number for the best factorized DD methods in each category in Table 1. In addition, we note the factorized DD methods with . Based on the performance comparison in the tables, we observe the followings:
@@ 6.2 Cross-Architecture Transferability
The standard benchmark in Table 1 only shows the performance of DD methods in terms of the specific ConvNet, while the distilled dataset is usually used to train other unseen network architectures. Therefore, measuring the DD performance on different architecture is also important for more comprehensive evaluation. However, there is not a standard set of architectures to evaluate cross-architecture transferability of DD algorithms, and the choice of architectures are highly different in the DD literature.
@@
In this survey, we collect the DD results on four popular architectures of ResNet-10/18/152 [41] and DenseNet-121 [46] to evaluate the cross-architecture transferability, as shown in Tables 3 and 4. By investigating these tables, we can obtain several observations as follows:
@@
  • There is a significant performance drop for nonfactorized DD methods when the distilled data are used to train unseen architectures (Table 3).
@@
  • The distilled data are not always possess the best performance across different architectures, which underlines the importance of evaluation on multiple architectures (Table 3).
@@
  • The factorized DD methods (IDC and KFS) possess better cross-architecture transferasbility i.e., suffer a smaller accuracy drop when encounter unseen architectures (Table 4).
@@ 7.1 Continual Learning
During to the training process, when there is a shift in the training data distribution, the model will suddenly lose its ability to predict the previous data distribution. This phenomenon is referred to as catastrophic forgetting and is common in deep learning. To overcome this problem, continual learning has been developed to incrementally learn new tasks while preserving performance on old tasks [94, 12]. A common method used in continual learning is the replay-based strategy, which allows a limited memory to store a few training examples for rehearsal in the following training. Therefore, the key to the replay-based strategy is to select highly informative training examples to store. Benefiting from extracting the essence of datasets, the dataset distillation technique has been employed to compress data for memory with limited storage [129, 11]. Because the incoming data have a changing distribution, the frequency is high for updating the elements in memory, which leads to strict requirements for the efficiency of dataset distillation algorithms. To conveniently embed dataset distillation into the replay-based strategy, Wiewel and Yang [116] and Sangermano et al. [99] decomposed the process of synthetic data generation by linear or nonlinear combination, and thus fewer parameters were optimized during dataset distillation. In addition to enhancing replay-based methods, dataset distillation can also learn a sequence of stable datasets, and the network trained on these stable datasets will not suffer from catastrophic forgetting [79].
@@ 7.2 Neural Architecture Search
7.3 Privacy Protection
As overparameterized neural networks easily memorize all the training data, there exists the risk for privacy leakage by inferring the well-trained networks [101, 73, 98, 44]. With a synthetic dataset, the network avoids explicitly training on the original dataset and consequently helps protect data privacy [132]. In addition, Chen et al. [18] employed gradient matching to distil private datasets by adding differential privacy noise to the gradients of synthetic data. Theoretically, Dong et al. [28] and Zheng and Li [130] built the connection between dataset distillation and differential privacy, and proved the superiority of dataset distillation in privacy preservation over conventional private data generation methods. However, Carlini et al. [10] argued that there exists flaws in both experimental and theoretical evidences in Dong et al. [28]: (1) membership inference baseline should include all training examples, while only $1\%$ of training points are used to match the synthetic sample size, which thus contributes to the high attack accuracy of baseline; and (2) the privacy analysis is based on a assumption stating that the output of learning algorithms follows a exponential distribution in terms of loss; and the assumption has already implied differential privacy.
@@ 7.4 Federated Learning
Federated learning has received increasing attention during training neural networks in the past few years due to its advantages in distributed training and private data protection [120]. The federated learning framework consists of multiple clients and one central server, and each client possesses exclusive data for training the corresponding local model [80]. In one round of federated learning, clients transmit the induced gradients or model parameters to the server after training with their exclusive data. Then the central server aggregates the received gradients or parameters to update the model parameters and broadcast new parameters to clients for the next round. In the federated learning scenario, the data distributed in different clients are often non-i.i.d data, which causes a biased minimum w.r.t. the local model and significantly hinders the convergence speed. Consequently, the data heterogeneity remarkably imposes a communication burden between the server and clients. Goetz and Tewari [34] proposed distilling a small set of synthetic data from original exclusive data by matching gradients, and then the synthetic data instead of a large number of gradients were transmitted to the server for model updating. Other distillation approaches such as BPTT [45, 131], KRR [103], and distribution matching [119], were also employed to compress the exclusive data to alleviate the communication cost at each transition round. However, Liu et al. [70] discovered that synthetic data generated by dataset distillation are still heterogeneous. To address the problem, these researchers proposed two strategies of dynamic weight assignment and meta knowledge sharing during the distillation process, which significantly accelerate the convergence speed of federated learning. Apart from compressing the local data, Pi et al. [90] also distilled data via trajectory matching in the server, which allows the synthetic data to possess global information. Then the distilled data can be used to fine-tune the server model for convergence speed up.
7.5 Adversarial Robustness
Deep learning networks with standard training are fragile with respect to adversarial attacks, and adding imperceptible perturbations to the input can completely fool the network [37]. Adversarial training has been widely used to address this problem by continuously feeding adversarial examples during the training process [78]. However, the generation of adversarial examples requires multistep gradient ascent, which considerably undermines training efficiency. Recently, Tsilivis et al. [111] and Wu et al. [117] employed dataset distillation to extract the information of adversarial examples and generate robust datasets. Then standard network training on the distilled robust dataset is sufficient to achieve satisfactory robustness to adversarial perturbations, which substantially saves computing resources compared to expensive adversarial training.
@@ 7.6 Other Applications
In addition to the continuous image domain, dataset distillation has also been extended to discrete data such as graphs [50, 49, 69] and recommender systems [96]. Jin et al. [50] first formulated this dataset distillation w.r.t. graph data, and successfully distilled a small, condensed graph via gradient matching. Due to the computational inefficiency in distilling pairwise relations for graph nodes, Jin et al. [49] turned to learning a probabilistic graph model, which allows a differentiable optimization w.r.t. the discrete graph. These authors also employed one-step strategy for further distillation speedup. Moreover, Liu et al. [71] accelerated the graph distillation through distribution matching from the perspective of receptive fields. For the recommender system, Sachdeva et al. [96] distilled a massive dataset to a continuous prior for tackling the discrete data problem.
@@
For backdoor attack, Liu et al. [72] considered to inject backdoor triggers into the small distilled dataset. Because the synthetic data has a small size, directly trigger insertion is less effective and also perceptible. Therefore, these coworkers turned to insert triggers into the target dataset during the dataset distillation procedure. In addition, they propose to iteratively optimize the triggers during the distillation process to preserve the triggers’ information for better backdoor attack performance.
8 Challenges and Directions
@@ 8.1 Challenges
For dataset distillation, there are two main ingredients: (1) measure the difference between the knowledge of synthetic data and real (original) data; and (2) update the synthetic data to narrow the knowledge difference. Based on these ingredients, we discuss the challenges of dataset distillation from the perspectives of (1) the definition of the difference in the knowledge between two types of datasets, (2) the form of synthetic datasets, (3) the evaluation of synthetic datasets, and (4) the theory behind dataset distillation.
Dataset distillation methods aim to transfer the knowledge of a target dataset into a small synthetic dataset to achieve a comparable generalization with respect to the original data distribution. Although the knowledge of datasets is an abstract notion, many DD approaches indirectly measure the difference in the knowledge base via various proxy losses. However, the existing definition w.r.t. the knowledge difference is inconsistent with the original comparable generalization. For example, the meta-loss $\mathcal{R}_{\mathcal{T}}(\texttt{alg}(\mathcal{S}))$ in the meta-learning framework measure the knowledge difference in the error in terms of a specific target dataset. Through this, one can only obtain a similar training error other than the test error w.r.t. the target dataset. As for the data matching framework, it intuitively makes knowledge concrete with a series of parameters or features, which are matched for knowledge transmission. Therefore, a more consistent definition of knowledge difference should be investigated to improve the dataset distillation performance.
Although most dataset distillation methods directly update the synthetic images, optimizing factorized codes and decoders that cooperatively generate synthetic data outperform the direct optimization by a large margin, as discussed in Section 6.1. From this, the predefined form of the code and decoder has a remarkable influence on the distillation performance. Many Factorization methods on synthetic data are based on intuition and lack rigorous analysis, and only some vague reasons are proposed, such as decreasing redundant information across different classes and learning more compact representations. Hence, it is necessary to explore the influence of the predefined factorization on the dataset distillation for learning a more informative distilled dataset.
In practical dataset distillation, the optimization of synthetic data is closely related to the specific network architecture (or kernels in KRR approach), and there is naturally a performance drop when the learned synthetic dataset is applied to other unseen architectures; please see Section 6.2. Therefore, it is one-sided to compare different DD methods only on one or a few architectures. To compare different DD methods in a more comprehensive scenario, it is essential to set a reasonable baseline consisting of multiple handpicked network architectures with different complexity.
@@
Although various DD approaches have been emerging, there have been few investigations that discuss the theory behind dataset distillation. Nevertheless, developing the theory is extremely necessary and will considerably promote the dataset distillation by directly improving the performance and also suggesting the correct direction of development. For example, the theory can help derive a better definition of the dataset knowledge and consequently increase the performance. In addition, the theoretical correlation between the distillation budget (the size of synthetic datasets) and performance can provide an upper bound on the test error, which can offer a holistic understanding and prevent researchers from blindly improving the DD performance. Hence, a solid theory is indispensable to take the development of DD to the next stage.

### v3

@@ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$
- $\mathbf{48.00}$ $\scriptscriptstyle\mathbf{\pm 0.3}$
TESLA [29] - - - - - -
$66.4$
$\scriptscriptstyle\pm 0.8$
- $\mathbf{48.9}$ $\scriptscriptstyle\mathbf{\pm 2.2}$ - - $\mathbf{34.8}$ $\scriptscriptstyle\mathbf{\pm 1.2}$ - - - - - -
@@
$53.7$
$\scriptscriptstyle\pm 0.7$
$45.07$
$\scriptscriptstyle\pm 0.5$
KIP [20]
$49.9$
$\scriptscriptstyle\pm 0.2$
@@
$\mathbf{16.6}$ $\scriptscriptstyle\mathbf{\pm 1.4}$ -
$15.9$
$\scriptscriptstyle\pm 1.1$
@@
$18.56$
$\scriptscriptstyle\pm 0.84$
@@
KFS [35] $\mathbf{59.8}$ $\scriptscriptstyle\mathbf{\pm 0.5}$ $\mathbf{47.0}$ $\scriptscriptstyle\mathbf{\pm 0.8}$ - - $\mathbf{49.5}$ $\scriptscriptstyle\mathbf{\pm 1.3}$ - $\mathbf{72.0}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{70.3}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ - - $\mathbf{71.4}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ - $\mathbf{75.0}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ $\mathbf{75.1}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ - - $\mathbf{76.3}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ -
@@ 6.1 Standard Benchmark
For a comprehensive comparison, we also present the test accuracy w.r.t. a random select search, coreset approach, and the original target dataset. For the coreset approach, the Herding algorithm [62] is employed in this survey due to its superior performance. Notably, for most of parametric DD methods, ConvNet [24, 63], which consists of three convolutional blocks, is the default architecture to distill the synthetic data; and all the test accuracy of distilled data are obtained by training the same architecture of ConvNet for a fair comparison.
The empirical comparison results associated with MNIST, FashionMNIST, SVHN, CIFAR-10/100, and Tiny ImageNet are presented in Table I. In spite of severe scale issues and few implementations on ImageNet-1K, we present the DD results of ImageNet-1K in Table II for completion. As shown in these tables, many methods are not tested on some datasets due to the lack of meaning for easy datasets or scalability problems for large datasets. We preserve these blanks in tables for a more fair and comprehensive comparison. To make the comparison clear, we use the orange number for the best of two nonfactorized DD methods and blue number for the best factorized DD methods in each category in Table I. In addition, we note the factorized DD methods with . Based on the performance comparison in the tables, we observe the followings:
@@ 6.2 Cross-Architecture Transferability
The standard benchmark in Table I only shows the performance of DD methods in terms of the specific ConvNet, while the distilled dataset is usually used to train other unseen network architectures. Therefore, measuring the DD performance on different architecture is also important for more comprehensive evaluation. However, there is not a standard set of architectures to evaluate cross-architecture transferability of DD algorithms, and the choice of architectures are highly different in the DD literature.
@@
In this survey, we collect synthetic data distilled by different methods on ConvNet to train other five popular architectures of ResNet-10/18/152 [2], DenseNet-121 [64], and Vision Transformer (ViT)  [^1] The ViT architecture has $4\times 4$ patch resolution, $6$ layers, $8$ heads, hidden dimension of $512$, and the MLP dimension of $512$. [65] to evaluate the cross-architecture transferability of distilled data, as shown in Tables III. By investigating these tables, we can obtain several observations as follows:
@@
  • There is a significant performance drop for nonfactorized DD methods when the distilled data are used to train unseen architectures.
@@
  • The distilled data are not always possess the best performance across different architectures, which underlines the importance of evaluation on multiple architectures.
@@
  • The factorized DD methods (IDC and KFS) possess better cross-architecture transferasbility i.e., suffer a smaller accuracy drop when encounter unseen architectures.
7 Data Modalities
Apart from image data that the majority of DD works focus on, some works leverage dataset distillation on multimodal data such as graph, text, etc.
7.1 Graph Data
Regarding a graph dataset of $\mathcal{T}=\{\mathbf{A},\mathbf{X},\mathbf{Y}\}$ with $N$ nodes, where $\mathbf{A}$, $\mathbf{X}$, and $\mathbf{Y}$ are the adjacency matrix, the node feature matrix, and the node labels, respectively, a graph neural network $\mathcal{G}$ is trained on $\mathcal{T}$ for node classification. Jin et al. [66] adopted the gradient matching framework to distill the graph $\mathcal{T}$ to a small synthetic graph $\mathcal{S}=\{\mathbf{A}^{\prime},\mathbf{X}^{\prime},\mathbf{Y}^{\prime}\}$ with $N^{\prime}$ nodes and $N^{\prime}\ll N$.
However, the number of parameters in the adjacency matrix $\mathbf{A}^{\prime}$ grows quadratically as $N^{\prime}$ increases, which greatly impedes the scalability of graph distillation. Considering the implicit correlation between the graph structure and node features, the authors employ a MLP-based model to predict the graph structure based on node features:
$\mathbf{A}_{ij}^{\prime}=\sigma\left(\frac{\operatorname{MLP}_{\Phi}\left(\left[\mathbf{x}_{i}^{\prime};\mathbf{x}_{j}^{\prime}\right]\right)+\operatorname{MLP}_{\Phi}\left(\left[\mathbf{x}_{j}^{\prime};\mathbf{x}_{i}^{\prime}\right]\right)}{2}\right),$ (39)
where $\sigma(\cdot)$ is the sigmoid function, $\text{MLP}_{\Phi}$ is parameterized with $\Phi$, and $\cdot$$\cdot$] denotes concatenation. With fixing $\mathbf{Y}^{\prime}$ to ease the difficulty of optimization, the loss function $\mathcal{L}(\mathbf{X}^{\prime},\Phi)$ of graph distillation for node classification is defined as
$\sum_{t=0}^{T-1}D\left(\nabla_{\bm{\theta}}\mathcal{R}\left(\mathcal{G}_{\bm{\theta}_{t}}\left({\Phi}\left(\mathbf{X}^{\prime}\right),\mathbf{X}^{\prime}\right),\mathbf{Y}^{\prime}\right),\nabla_{\bm{\theta}}\mathcal{R}\left(\mathcal{G}_{\bm{\theta}_{t}}(\mathbf{A},\mathbf{X}),\mathbf{Y}\right)\right).$ (40)
Moreover, to bypass second-order meta-gradient computation, Liu et al. [30] accelerated the graph distillation through distribution matching from the perspective of receptive fields, which present the local graph w.r.t. a specific node.
However, for the graph classification problem where adjacency matrix is discrete, the conventional DD algorithms are no longer applicable. To distill synthetic graphs with discrete adjacency matrix $\mathbf{A}^{\prime}\in\{0,1\}^{|\mathbf{A}^{\prime}|}$, Jin et al. [67] formulate $\mathbf{A}^{\prime}$ as a probabilistic graph with Bernoulli distribution
$\mathbf{P}_{\Omega_{ij}}\left(\mathbf{A}_{ij}^{\prime}\right)=\mathbf{A}_{ij}^{\prime}\sigma\left(\Omega_{ij}\right)+\left(1-\mathbf{A}_{ij}^{\prime}\right)\sigma\left(-\Omega_{ij}\right),$ (41)
where $\sigma$ is the sigmoid function and $\Omega_{ij}\in\mathbb{R}$ is the learnable parameter. Then with the reparameterization trick, the parameter $\Omega$ becomes differentiable and $\mathbf{A}^{\prime}$ is accordingly synthesized by optimizing $\Omega$.
7.2 Text Data
For the discrete text data, Li and Li [68] converted each piece of text into a continuous embedding matrix to adjust DD algorithms, and then a few high-informative synthetic embedding matrices were distilled for text classification. To fine-tune large language models like BERT [69] with the distilled text data, Maekawa et al. [70] further optimized the soft label by decreasing the KL divergence between the attention probabilities of distilled data and the model during the DD process, which facilitates efficient knowledge transfer to transformer models.
@@ 8.1 Continual Learning
During to the training process, when there is a shift in the training data distribution, the model will suddenly lose its ability to predict the previous data distribution. This phenomenon is referred to as catastrophic forgetting and is common in deep learning. To overcome this problem, continual learning has been developed to incrementally learn new tasks while preserving performance on old tasks [62, 71]. A common method used in continual learning is the replay-based strategy, which allows a limited memory to store a few training examples for rehearsal in the following training. Therefore, the key to the replay-based strategy is to select highly informative training examples to store. Benefiting from extracting the essence of datasets, the dataset distillation technique has been employed to compress data for memory with limited storage [24, 72]. Because the incoming data have a changing distribution, the frequency is high for updating the elements in memory, which leads to strict requirements for the efficiency of dataset distillation algorithms. To conveniently embed dataset distillation into the replay-based strategy, Wiewel and Yang [73] and Sangermano et al. [74] decomposed the process of synthetic data generation by linear or nonlinear combination, and thus fewer parameters were optimized during dataset distillation. Gu et al. [75] develop the dynamic memory to maintain both distilled and real data, which helps better summarize data information with the distilled data.
In addition to enhancing replay-based methods, dataset distillation can also learn a sequence of stable datasets, and the network trained on these stable datasets will not suffer from catastrophic forgetting [76].
@@ 8.3 Federated Learning
Federated learning has received increasing attention during training neural networks in the past few years due to its advantages in distributed training and private data protection [79]. The federated learning framework consists of multiple clients and one central server, and each client possesses exclusive data for training the corresponding local model [80]. In one round of federated learning, clients transmit the induced gradients or model parameters to the server after training with their exclusive data. Then the central server aggregates the received gradients or parameters to update the model parameters and broadcast new parameters to clients for the next round. In the federated learning scenario, the data distributed in different clients are often non-i.i.d data, which causes a biased minimum w.r.t. the local model and significantly hinders the convergence speed. Consequently, the data heterogeneity remarkably imposes a communication burden between the server and clients. Goetz and Tewari [81], Wang et al. [82] proposed distilling a small set of synthetic data from original exclusive data by matching gradients, and then the synthetic data instead of a large number of gradients were transmitted to the server for model updating. Other distillation approaches such as BPTT [83, 84], KRR [85], and distribution matching [86], were also employed to compress the exclusive data to alleviate the communication cost at each transition round. However, Liu et al. [87] discovered that synthetic data generated by dataset distillation are still heterogeneous. To address the problem, these researchers proposed two strategies of dynamic weight assignment and meta knowledge sharing during the distillation process, which significantly accelerate the convergence speed of federated learning. Apart from compressing the local data, Pi et al. [88] also distilled data via trajectory matching in the server, which allows the synthetic data to possess global information. Then the distilled data can be used to fine-tune the server model for convergence speed up.
@@ 8.4 Other Applications
9 Challenges and Future Directions
@@ 9.1 Challenges
Scalability. Although many approaches have been proposed to decrease memory utilization [29] and accelerate convergence [55] for efficient DD, the scalability of DD is still a serious issue due to the hardness of optimizing numerous hyperparameters (e.g., pixels for image data). For example, when distilling the $1000$-class ImageNet with the input resolution of $224\times 224$ and IPC of $50$, totally $224\times 224\times 1000\times 50\approx 2.5\text{billon}$ hyperparameters (pixels) are required to simultaneously optimize. Therefore, how to effectively search a feasible solution over the extremely high-dimensional hyperparameter space is important to scale DD algorithms to more complicated data (e.g., image data with higher resolution, 3D data).
Multimodal dataset distillation. Whereas DD algorithms have achieved remarkable progress in terms of image [11, 24], graph [66], text [68], tabular [94], and recommendation system data [95], their applications on other data modalities are still underexplored. Especially for the modalities such as video that rich spatial and temporal information are tangled with each other in a single data [96], it is still challenging for DD algorithms to simultaneously distill the spatial and temporal features into a small volume of data.
DD with complex labels. The majority of DD techniques lie in the single-label classification task where its label space is an integer set. As for other challenging tasks such as detection [97] or segmentation [98], their label spaces locate in much higher dimensional spaces. Unlike DD with single-label classification that the specific number of synthetic images (IPC) are pre-assigned for each class, enumerated assignment is impractical when the label space is highly-dimensional. Therefore, to represent the rich label space, high-dimensional labels of synthetic data require cautious adjusting, which significantly hinders the DD efficiency especially when jointly optimized with the input data.
@@
Theory of dataset distillation. Although various DD algorithms have been emerging, there have been few investigations on discussing the theory behind dataset distillation. Nevertheless, developing the theory is extremely necessary and will considerably promote the dataset distillation by directly improving the performance and also suggesting the correct direction of development. For example, the theory can help derive a better definition of the dataset knowledge and consequently increase the performance. In addition, the theoretical correlation between the distillation budget (the size of synthetic datasets) and performance can provide an upper bound on the test error, which can offer a holistic understanding and prevent researchers from blindly improving the DD performance. Hence, a solid theory is indispensable to take the development of DD to the next stage.